OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (anonymous):

like how know 1 replies when i ask o_O

OpenStudy (anonymous):

\[f(2)=-3f(1)-2f(0)=-3(12)-2(-4)=-28\] It's gotta be the first choice. The second one doesn't have the right term for \(f(1)\).

OpenStudy (anonymous):

ty do u know y like i dont get it at all

OpenStudy (anonymous):

:?